Find all solutions $x$ to the equation
5(2x + 1) + 5*2 = 5*x + 5(x + 3).

Answer by ikleyn(52936)

.
Find all solutions $x$ to the equation
5(2x + 1) + 5*2 = 5*x + 5(x + 3).
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Simplify, step by step

    5(2x + 1) + 5*2 = 5*x + 5(x + 3)

     10x + 5 + 10   = 5x + 5x + 15

     10x + 15       = 10x + 15.


At this point, you see that the left side is identical to the right side.


It means that there are infinitely many solutions: every real number x is the solution.

Solved, with comprehensive explanations.